OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation. 3a/7 + 2/3 = 7/9

OpenStudy (kc_kennylau):

We have LCM and GCD, I think you mean LCM: LCM(7,3,9)=LCM(7,9)=7*9=63